# arXiv:2602.23342v1[cs.DB] 26 Feb 2026

### v1

@@ 1. Introduction
Interestingly, as shown in Figures 1(a), (b) and (c), we observed that the CPU costs contributed to a large proportion of the query latency in all existing on-disk graph-based index systems (i.e., DiskANN, Starling, and PipeANN). We further systematically validate the above observation with an adapted roofline model. Contrary to current wisdom, i.e., the performance of on-disk graph-based index systems is limited by I/O costs, our analyzed results from the roofline model reveal the performance of these systems changes from I/O-bound to compute-bound with the rising of the vector dimensionality. For example, the performance of DiskANN, Starling and PipeANN is actually I/O-bound on 96-dimensional DEEP10M, but it turned to compute-bound on 960-dimensional GIST1M, we will elaborate the details shortly in Section 2.2.
@@ 2. Preliminaries and Analysis
Let $\mathcal{D}$ be a dataset of $d$-dimensional vectors. For a query vector $q$, vector similarity search identifies $k$ vectors in $\mathcal{D}$ with the smallest Euclidean distances to $q$. In this work, we design an on-disk, graph-based index system that enables efficient ANNS on large-scale high-dimensional vector dataset $\mathcal{D}$ with a limited memory budget $B$. Specifically, the size of vector dataset $\mathcal{D}$ vastly exceeds the available memory budget $B$. The objective of this work is to maximize search performance (i.e., high throughput and low latency), while maintaining good result accuracy (i.e., high recall) for high-dimensional vectors (e.g., those with hundreds or thousands of dimensions). In this section, we first present the general idea of existing on-disk graph-based index systems for high-dimensional vector similarity search in Section 2.1, then identify the performance bottleneck of these systems on two vector datasets with different dimensionalities via adapted roofline model in Section 2.2.
@@ 2.1. Existing On-disk Graph-based Indexes
Distance computation. For a given query $q$, DiskANN generates a lookup table (LUT) in memory to store the distances of query sub-vector and the cluster centroids of every data sub-vector space, see the in-memory LUT in Figure 3. The LUT can be used to obtain the approximate distance between query vector and candidate data vectors and guide the best vectors (and neighbors) to read from disk. For example, the approximate distance between query $q$ and data vector $2$ can be computed by summing the values which are obtained by accessing 1 to 4 of LUT. The approximate distance can be refined by computing the exact distance between the query vector and full precision data vectors, and the final ANNS results can be returned by re-ranking these candidates with exact distances.
Beam search strategy. To reduce the number of I/Os, DiskANN devises beam search strategy, which fetches the neighborhoods of a small number of the closest points (a.k.a., beam width $BW$) in one shot as reading a few random sectors from an SSD takes almost the same amount of time as reading one sector.
@@
2.2. Roofline Model Analysis
@@
Achieved performance evaluation. We measure the achieved performance of these three on-disk graph-based index systems on two vector datasets with the used machine. We use all 48 available threads to maximize the system utilization. As depicted in Figure 4, the achieved performance of DiskANN, Starling and PipeANN on 96-dimensional DEEP10M are 3.96 GFLOP/s, 11.37 GFLOP/s and 5.09 GFLOP/s, respectively. On DEEP10M, the performance of DiskANN and Starling is limited by the actual SSD bandwidth as it only achieved 6.5 GB/s, which is far below its peak bandwidth $\beta$ = 11.1 GB/s. The reason is that the beam search width $BW$ is configured as 8, which limits the depth of I/O queue in them (Haas and Leis, 2023). PipeANN achieves a higher performance than DiskANN by employing a deeper I/O queue (i.e., 32) in its pipeline, which utilizes the available SSD bandwidth better. On 960-dimensional GIST1M, the achieved performance of DiskANN, Starling and PipeANN are 18.94 GFLOP/s, 18.85 GFLOP/s, and 18.28 GFLOP/s, respectively. All of these solutions are significantly below the peak performance $\pi$ = 33.6 GFLOP/s. The reason is that the peak performance is derived at the ideal situation, i.e., it only takes 3 cycles for each FLOP as it assumes all the data are in L1 cache. However, the PQ compressed vector of each data vector in 960-dimensional GIST1M is impossible to entirely store on L1 cache, which means it takes more CPU cycles for each FLOP then the peak performance $\pi$ will be smaller than the maximum 33.6 GFLOP/s in real-world systems.
@@ 3. Index Layout of AlayaLaser
As confirmed by the adapted roofline model in Section 2, the performance bottleneck of existing on-disk graph-based index systems is the expensive computational costs. In this section, we devise an effective index layout in AlayaLaser to overcome it. However, it is not trivial to achieve it due to the technical challenges in Section 3.1. We next design the SIMD-friendly layout of AlayaLaser in Section 3.2, and present the node caching scheme of AlayaLaser in Section 3.3.
@@ 3.1. Technical Challenges
C1: Limitation of sequential computation pattern. As discussed in Section 2.2, the computation cost of existing on-disk graph-based index systems (i.e., DiskANN, Starling, and PipeANN) is dominated by the approximate distance calculation from the quantization codewords. In particular, it incurs numerous LUT lookups. Taking Figure 3 as an example, it consumes 12 LUT lookups to derive the approximate distances from the query $q$ to the neighbors of data vector $1$. Although LUT is typically small enough to fit within the L1 cache of CPU, which allows low-latency lookups, the lookup process must proceed step by step, with each stage dependent on the completion of the prior one. For example, as shown in Figure 3, 0.9, 2.8, 4.8 and 0.3 are accessed by looking up the LUT step by step then used to estimate the approximate distance between query vector $q$ and data vector $2$. Thus, the above computation procedure cannot be effectively parallelized and executed by exploiting SIMD instructions in modern CPUs due to the irregular memory access and inefficient vectorization. Thus, it poses the first technical challenge to overcome the compute-bound of existing on-disk graph-based index systems.
@@ 3.2. SIMD-friendly Layout of AlayaLaser
Design of SIMD-friendly layout in AlayaLaser. The overview of the SIMD-friendly layout in AlayaLaser is illustrated at the bottom of Figure 5, we introduce the designs of it as follows. The key insight behind the design is that we observed that the quantization scheme (e.g., PQ) treats all dimensions in data vector uniformly and allocates equivalent bits to different dimensions, which results in suboptimal disk page utilization. To overcome this, we first utilize Principal Component Analysis (PCA) to project the original high-dimensional data vectors into a low-dimensional subspace that captures the majority of the data variance. In particular, the Principal Component Analysis (PCA) transformer will generate two parts for each data vector $i$: a principal component $V_{i}^{d}$, and a residual component $V_{i}^{r}$, the exact distance between the query $q$ and data vector $i$ is preserved when the residual component is also included (Bishop and Nasrabadi, 2006; Jégou et al., 2010; Babenko and Lempitsky, 2014). Second, we only apply a high-ratio quantizer (e.g., Product Quantization (PQ) (Jegou et al., 2010) or RaBitQ (Gao and Long, 2024)) on the principal component of the neighboring data vectors. For example, the neighboring principal components of data vector $1$ are $V_{2}^{d}$, $V_{6}^{d}$ and $V_{8}^{d}$. They are compressed via a high-ratio quantizer (Gao and Long, 2024). We obtain the highly compact quantized codewords, as shown in Figure 5. To minimize the quantization error, we apply an orthogonal rotation to balance the variance distribution across the dimensions of principal components, which allows the quantizer to achieve higher precision for a given code size. To fully exploit the SIMD instructions on CPU, we re-arrange these quantized codewords in an interleaved manner. In particular, we decompose the quantized codeword of each data vector into individual sub-codes, then sequentially store the same position sub-codes of different neighbors. As illustrated in Figure 5, the sub-codes of “a” and “b” from the neighboring data vectors 2, 6, 8 are stored adjacently after the rearrangement transformation. Last, the compact and rearranged quantized codewords of the neighbors are appended.
@@
Determining the graph out-degree $R$: The graph out-degree $R$ directly determines the disk page size in AlayaLaser. The above SIMD-friendly layout of AlayaLaser enables the using of the commonly-used out-degrees (e.g., $R=64$ or $96$) in existing on-disk graph-based index systems (e.g., DiskANN) while keeping small page size. The core reason is that AlayaLaser only stores the quantization codewords of the principal component $V_{i}^{d}$, instead of the original data vector. For example, on 960-dimensional GIST1M, the page size of AlayaLaser is 8KB for out-degree $R=64$ and it also is only 12KB if the out-degree is $R=128$. It confirms the design of our SIMD-friendly layout ensures the high I/O performance (i.e., IOPS). In addition, the performance gains diminish rapidly with the increase of graph out-degree $R$ (Fu et al., 2017). Thus, the out-degree $R$ does not exceed 128 in AlayaLaser.
@@
Running example: The example in Figure 6 illustrates the distance computation procedure with SIMD-friendly layout. In particular, it loads the quantized codewords of the same position from different neighbors into a SIMD register. For example, it first loads the sub-codes “a” from neighboring data vectors 2, 6, 8 of data vector 1. Then, the corresponding LUT for the sub-codes “a” is loaded into another SIMD register. After that, the distances between the each sub-code of the query and vectors 2, 6, 8 are obtained via applying SIMD shuffle instruction (Guide, 2011) on these two SIMD registers, which are 0.9, 4.8 and 0.9, respectively, as depicted by colors in LUT “a”. Then applying the same procedure on the sub-codes “b”, and compute the approximate distance between query vector and data vectors $2$, 66 and $8$. Compared to DiskANN, AlayaLaser transforms the strict serial LUT accesses into a parallel execution procedure via the SIMD-friendly layout, which significantly reduces the computational cost. In particular, as shown in Figure 3, DiskANN computes the approximate distances among the neighboring data vectors $2$, 66 and $8$ of data vector $1$ with query vector $q$ by invoking 24 scalar loads and 12 adds (Blalock and Guttag, 2017). However, our AlayaLaser computes these approximate distances by 4 SIMD loads, 2 SIMD shuffles and 2 SIMD adds. It is worth pointing out that AlayaLaser works on the quantized codewords of principal component of data vector, which is smaller than the PQ compressed vectors in DiskANN. Thus, the computation cost of AlayaLaser will be further reduced.
3.3. Node Caching Scheme in AlayaLaser
@@ 4.2. Adaptive Beam Expansion Strategy
AlayaLaser follows the beam search strategy of DiskANN to reduce the I/O times during graph traversal. The core of beam search is exploring a batch of nodes at each hop, and the batch size is determined by the beam width $BW$. Generally, a larger beam width $BW$ increases the likelihood to find a shorter search path, but incurs higher computational and I/O overhead. Existing on-disk graph-based index systems (i.e., DiskANN and Starling) typically configure a fixed beam width $BW$ throughout the entire search process, which is inefficient as it cannot adapt to the dynamic context of the search. Specifically, in the early stage of search processing, the query vector is far away from the visited nodes in the graph-based index, the fixed wide beam width may waste cost as it visits many unqualified nodes, i.e., which are not be the final result set, by “over expansion”, see the gray dashed arrows in Figure 8(b). However, in the final stage of search processing, the fixed beam width $BW$ may lead to insufficient exploration as the search procedure enters a dense region of candidates, which leads to long search path, consequently, results to higher search latency. Inspired by (Peng et al., 2023), we devise an adaptive beam expansion strategy in AlayaLaser to overcome this issue. In particular, the beam width $BW$ at the i-th hop is determined by $min(2^{i},BW_{D})$, where $BW_{D}$ is the default beam width and it is 2-4 times larger than that of DiskANN. The core idea of this method is enabling coarse-to-fine search: a narrow initial beam reduces wasteful I/Os in distant regions, while the exponential growth ($2^{i}$) enables a rapid transition to intensive exploration when the search converges to the targets, see the red arrows in Figure 8(b).

### v2

# AlayaLaser: Efficient Index Layout and Search Strategy for Large-scale High-dimensional Vector Similarity Search

@@ 1. Introduction
Interestingly, as shown in Figures 1(a), (b) and (c), we observed that the CPU costs contributed to a large proportion of the query latency in all existing on-disk graph-based index systems (i.e., DiskANN, Starling, and PipeANN). We further systematically validate the above observation with an adapted roofline model. Contrary to current wisdom, i.e., the performance of on-disk graph-based index systems is limited by I/O costs, our analyzed results from the roofline model reveal the performance of these systems changes from I/O-bound to compute-bound with the rising of the vector dimensionality. For example, the performance of DiskANN, Starling and PipeANN is actually I/O-bound on 96-dimensional DEEP10M, but it turned to compute-bound on 960-dimensional GIST1M, we will elaborate the details shortly in Section $2.$
@@ 2. Preliminaries and Analysis
Let $\mathcal{D}$ be a dataset of $d$-dimensional vectors. For a query vector $q$, vector similarity search identifies $k$ vectors in $\mathcal{D}$ with the smallest Euclidean distances to $q$. In this work, we design an on-disk, graph-based index system that enables efficient ANNS on large-scale high-dimensional vector dataset $\mathcal{D}$ with a limited memory budget $B$. Specifically, the size of vector dataset $\mathcal{D}$ vastly exceeds the available memory budget $B$. The objective of this work is to maximize search performance (i.e., high throughput and low latency), while maintaining good result accuracy (i.e., high recall) for high-dimensional vectors (e.g., those with hundreds or thousands of dimensions). In this section, we first present the general idea of existing on-disk graph-based index systems for high-dimensional vector similarity search in Section 2.1, then identify the performance bottleneck of these systems on two vector datasets with different dimensionalities via adapted roofline model in Section $2.$
@@ 2.1. Existing On-disk Graph-based Indexes
Distance computation. For a given query $q$, DiskANN generates a lookup table (LUT) in memory to store the distances of query sub-vector and the cluster centroids of every data sub-vector space, see the in-memory LUT in Figure 3. The LUT can be used to obtain the approximate distance between query vector and candidate data vectors and guide the best vectors (and neighbors) to read from disk. For example, the approximate distance between query $q$ and data vector 22 can be computed by summing the values which are obtained by accessing 1 to 4 of LUT. The approximate distance can be refined by computing the exact distance between the query vector and full precision data vectors, and the final ANNS results can be returned by re-ranking these candidates with exact distances.
Beam search strategy. To reduce the number of I/Os, DiskANN devises beam search strategy, which fetches the neighborhoods of a small number of the closest points (a.k.a., beam width BWBW) in one shot as reading a few random sectors from an SSD takes almost the same amount of time as reading one sector.
@@
$2.$ Roofline Model Analysis
@@
Achieved performance evaluation. We measure the achieved performance of these three on-disk graph-based index systems on two vector datasets with the used machine. We use all 48 available threads to maximize the system utilization. As depicted in Figure 4, the achieved performance of DiskANN, Starling and PipeANN on 96-dimensional DEEP10M are 3.96 GFLOP/s, 11.37 GFLOP/s and 5.09 GFLOP/s, respectively. On DEEP10M, the performance of DiskANN and Starling is limited by the actual SSD bandwidth as it only achieved 6.5 GB/s, which is far below its peak bandwidth $\beta$ = 11.1 GB/s. The reason is that the beam search width BWBW is configured as 8, which limits the depth of I/O queue in them (Haas and Leis, 2023). PipeANN achieves a higher performance than DiskANN by employing a deeper I/O queue (i.e., 32) in its pipeline, which utilizes the available SSD bandwidth better. On 960-dimensional GIST1M, the achieved performance of DiskANN, Starling and PipeANN are 18.94 GFLOP/s, 18.85 GFLOP/s, and 18.28 GFLOP/s, respectively. All of these solutions are significantly below the peak performance $\pi$ = 33.6 GFLOP/s. The reason is that the peak performance is derived at the ideal situation, i.e., it only takes 3 cycles for each FLOP as it assumes all the data are in L1 cache. However, the PQ compressed vector of each data vector in 960-dimensional GIST1M is impossible to entirely store on L1 cache, which means it takes more CPU cycles for each FLOP then the peak performance $\pi$ will be smaller than the maximum 33.6 GFLOP/s in real-world systems.
@@ 3. Index Layout of AlayaLaser
As confirmed by the adapted roofline model in Section 2, the performance bottleneck of existing on-disk graph-based index systems is the expensive computational costs. In this section, we devise an effective index layout in AlayaLaser to overcome it. However, it is not trivial to achieve it due to the technical challenges in Section 3.1. We next design the SIMD-friendly layout of AlayaLaser in Section 3.2, and present the node caching scheme of AlayaLaser in Section $3.$
@@ 3.1. Technical Challenges
C1: Limitation of sequential computation pattern. As discussed in Section 2.2, the computation cost of existing on-disk graph-based index systems (i.e., DiskANN, Starling, and PipeANN) is dominated by the approximate distance calculation from the quantization codewords. In particular, it incurs numerous LUT lookups. Taking Figure 3 as an example, it consumes 12 LUT lookups to derive the approximate distances from the query $q$ to the neighbors of data vector 11. Although LUT is typically small enough to fit within the L1 cache of CPU, which allows low-latency lookups, the lookup process must proceed step by step, with each stage dependent on the completion of the prior one. For example, as shown in Figure 3, 0.9, 2.8, 4.8 and 0.3 are accessed by looking up the LUT step by step then used to estimate the approximate distance between query vector $q$ and data vector 22. Thus, the above computation procedure cannot be effectively parallelized and executed by exploiting SIMD instructions in modern CPUs due to the irregular memory access and inefficient vectorization. Thus, it poses the first technical challenge to overcome the compute-bound of existing on-disk graph-based index systems.
@@ 3.2. SIMD-friendly Layout of AlayaLaser
Design of SIMD-friendly layout in AlayaLaser. The overview of the SIMD-friendly layout in AlayaLaser is illustrated at the bottom of Figure 5, we introduce the designs of it as follows. The key insight behind the design is that we observed that the quantization scheme (e.g., PQ) treats all dimensions in data vector uniformly and allocates equivalent bits to different dimensions, which results in suboptimal disk page utilization. To overcome this, we first utilize Principal Component Analysis (PCA) to project the original high-dimensional data vectors into a low-dimensional subspace that captures the majority of the data variance. In particular, the Principal Component Analysis (PCA) transformer will generate two parts for each data vector $i$: a principal component $V_{i}^{d}$, and a residual component $V_{i}^{r}$, the exact distance between the query $q$ and data vector $i$ is preserved when the residual component is also included (Bishop and Nasrabadi, 2006; Jégou et al., 2010; Babenko and Lempitsky, 2014). Second, we only apply a high-ratio quantizer (e.g., Product Quantization (PQ) (Jegou et al., 2010) or RaBitQ (Gao and Long, 2024)) on the principal component of the neighboring data vectors. For example, the neighboring principal components of data vector 11 are $V_{2}^{d}$, $V_{6}^{d}$ and $V_{8}^{d}$. They are compressed via a high-ratio quantizer (Gao and Long, 2024). We obtain the highly compact quantized codewords, as shown in Figure 5. To minimize the quantization error, we apply an orthogonal rotation to balance the variance distribution across the dimensions of principal components, which allows the quantizer to achieve higher precision for a given code size. To fully exploit the SIMD instructions on CPU, we re-arrange these quantized codewords in an interleaved manner. In particular, we decompose the quantized codeword of each data vector into individual sub-codes, then sequentially store the same position sub-codes of different neighbors. As illustrated in Figure 5, the sub-codes of “a” and “b” from the neighboring data vectors 2, 6, 8 are stored adjacently after the rearrangement transformation. Last, the compact and rearranged quantized codewords of the neighbors are appended.
@@
Determining the graph out-degree $R$: The graph out-degree $R$ directly determines the disk page size in AlayaLaser. The above SIMD-friendly layout of AlayaLaser enables the using of the commonly-used out-degrees (e.g., $R=64$ or 9696) in existing on-disk graph-based index systems (e.g., DiskANN) while keeping small page size. The core reason is that AlayaLaser only stores the quantization codewords of the principal component $V_{i}^{d}$, instead of the original data vector. For example, on 960-dimensional GIST1M, the page size of AlayaLaser is 8KB for out-degree $R=64$ and it also is only 12KB if the out-degree is $R=128$. It confirms the design of our SIMD-friendly layout ensures the high I/O performance (i.e., IOPS). In addition, the performance gains diminish rapidly with the increase of graph out-degree $R$ (Fu et al., 2017). Thus, the out-degree $R$ does not exceed 128 in AlayaLaser.
@@
Running example: The example in Figure 6 illustrates the distance computation procedure with SIMD-friendly layout. In particular, it loads the quantized codewords of the same position from different neighbors into a SIMD register. For example, it first loads the sub-codes “a” from neighboring data vectors 2, 6, 8 of data vector 1. Then, the corresponding LUT for the sub-codes “a” is loaded into another SIMD register. After that, the distances between the each sub-code of the query and vectors 2, 6, 8 are obtained via applying SIMD shuffle instruction (Guide, 2011) on these two SIMD registers, which are 0.9, 4.8 and 0.9, respectively, as depicted by colors in LUT “a”. Then applying the same procedure on the sub-codes “b”, and compute the approximate distance between query vector and data vectors 22, 66 and 88. Compared to DiskANN, AlayaLaser transforms the strict serial LUT accesses into a parallel execution procedure via the SIMD-friendly layout, which significantly reduces the computational cost. In particular, as shown in Figure 3, DiskANN computes the approximate distances among the neighboring data vectors 22, 66 and 88 of data vector 11 with query vector $q$ by invoking 24 scalar loads and 12 adds (Blalock and Guttag, 2017). However, our AlayaLaser computes these approximate distances by 4 SIMD loads, 2 SIMD shuffles and 2 SIMD adds. It is worth pointing out that AlayaLaser works on the quantized codewords of principal component of data vector, which is smaller than the PQ compressed vectors in DiskANN. Thus, the computation cost of AlayaLaser will be further reduced.
$3.$ Node Caching Scheme in AlayaLaser
@@ 4.2. Adaptive Beam Expansion Strategy
AlayaLaser follows the beam search strategy of DiskANN to reduce the I/O times during graph traversal. The core of beam search is exploring a batch of nodes at each hop, and the batch size is determined by the beam width BWBW. Generally, a larger beam width BWBW increases the likelihood to find a shorter search path, but incurs higher computational and I/O overhead. Existing on-disk graph-based index systems (i.e., DiskANN and Starling) typically configure a fixed beam width BWBW throughout the entire search process, which is inefficient as it cannot adapt to the dynamic context of the search. Specifically, in the early stage of search processing, the query vector is far away from the visited nodes in the graph-based index, the fixed wide beam width may waste cost as it visits many unqualified nodes, i.e., which are not be the final result set, by “over expansion”, see the gray dashed arrows in Figure 8(b). However, in the final stage of search processing, the fixed beam width BWBW may lead to insufficient exploration as the search procedure enters a dense region of candidates, which leads to long search path, consequently, results to higher search latency. Inspired by (Peng et al., 2023), we devise an adaptive beam expansion strategy in AlayaLaser to overcome this issue. In particular, the beam width BWBW at the i-th hop is determined by $min(2^{i},BW_{D})$, where $BW_{D}$ is the default beam width and it is 2-4 times larger than that of DiskANN. The core idea of this method is enabling coarse-to-fine search: a narrow initial beam reduces wasteful I/Os in distant regions, while the exponential growth ($2^{i}$) enables a rapid transition to intensive exploration when the search converges to the targets, see the red arrows in Figure 8(b).